# Enhancing Trust in LLM-Generated Code Summaries with Calibrated Confidence Scores

### v2

@@ 6.1. Limitation: Summary Intent Capture
The notion of correctness in our work is based on similarity to human-produced summaries. However, human developers have different intents (how/what/why) when writing summaries (Hu et al., 2022; Geng et al., 2024). These different intents are not captured by the similarity metrics. A reliable way of integrating intent into summary evaluation metrics remains for future work.
That said, it is important to note that prior work indicates that human-produced summaries are inherently valuable (Stapleton et al., 2020) and that the measure we used is an excellent proxy of human perceptions of similarity to human summaries (Haque et al., 2022; Mastropaolo et al., 2024).
6.2. Limitation: Experimental Design
In this paper, we primarily experimented with two programming languages, two prompting techniques, three models, and primarily with one similarity metric. Our findings in this paper are limited to these components. Prompting techniques can vary greatly, and the findings may differ significantly based on the prompting technique, as well as the programming language and metric used. Addressing all the languages, models, and prompting techniques is beyond the scope of this paper. However, Java and Python are two very different programming languages, and the models we used exhibit diversity in size and training. We believe that our findings will be well-generalized beyond these languages, models, and other parameters.
@@ 6.3. Limitation: Platt scaling
Platt scaling is a popular and widely used rescaling strategy. However, another very popular rescaling technique is temperature scaling, which is found to be better than Platt scaling in certain scenarios. In this paper, we could not use temperature scaling because it requires access to complete softmax layers, which are unavailable from the OpenAI [^2] https://openai.com and TogetherAI [^3] https://www.together.ai APIs. In the future, we hope to incorporate temperature scaling into our experiments when we have access to the complete softmax layer.
@@ 6.4. Self-Reflection Calibration
By “reflective” we mean the ability of an instruction-tuned model to evaluate its own output. We used two reflective measures to compute model confidence. One is logit-based, and the other one is verbalized. In the logit-based measure, we present the model with the code and model-generated summary, and ask the model to rate the quality of the summary by generating a true/false output (we used similar prompts to those proposed by Spiess et al. (Spiess et al., 2025)). We review the top five tokens, by their probabilities, and select the probability associated with the “True” token. For the verbalized measure, we tried both zero-shot and few-shot learning approaches and asked the model to assign a score from 1 to 4 based on similarity, following the approach used by Haque et al. (Haque et al., 2022)’s tool for human validation. In regards to the reflective logit, when we prompt the model to generate true/false responses, CodeLlama-70b produced invalid responses on extensive prompt variations since it is not instruction-tuned. Additionally, we are interested in the logit associated with the token “True”. If the model does not generate “True”, it is impossible to retrieve the “True” logit unless we have access to the top-k log-probability tokens and “True” is one of them. Among our models, we only have this access for GPT-3.5-Turbo due to API limitations. Consequently, we only evaluated reflective logits on the GPT-3.5-Turbo model. However, the model was highly over-confident with 94% of its confidences over 0.9, achieving a negative skill score. Rescaling transformed all the logits to near the unskilled, base rate.
We also attempted zero-shot and few-shot verbalized confidence measures by prompting the model to assign either 1, 2, 3, or 4 to score the similarity between the reference and the generated output. This design choice is inspired by human evaluations by Haque et al. (Haque et al., 2022). In this setup, CodeLlama-70b failed to generate any score, but in the few-shot learning scenario, CodeLlama-70b is able to generate scores. During few-shot learning, the shots were taken from the dataset proposed by Haque et al. For both zero-shot and few-shot (4 shots) settings, we found that the models were not well-calibrated. Note that we tried a higher temperature and generated multiple samples. Majority voting from multiple samples did not change the calibration outcome.
@@ 7. Related Work
8. Conclusion
Code summaries are valuable to maintainers, and LLMs can generate good summaries from code. However, LLM-generated summary quality is very variable: sometimes similar to human-written summaries, and sometimes not. This indicates a need for developers to know when a summary might be sufficiently similar to a human-written summary. This is a calibration problem: can the model provide a reliable indication of its confidence that a generated summary is “good enough”? We address this problem by selecting a metric known from the literature to be a good proxy for similarity to human-produced summaries, and then thresholding this metric to get an indication of sufficient similarity. We then evaluate the calibration of a range of prompting approaches, together with Platt scaling. We find that in some cases, the first few tokens may provide the best calibration; we also find both retrieval-augmented few-shot prompting and ASAP can be advantageous. We also find that reflection-based approaches do not provide better calibration. To our knowledge, we are the first to evaluate whether LLM-based approaches to code summarization are well calibrated; our work contributes both a methodology, and several useful approaches, to this important question.
@@ 9. Data Availability
Our code and dataset are made available at 10.5281/zenodo.10858303.

### v3

@@ 6. Discussion: Threats & Limitations
6.1. Do LLMs surpass code summaries found in open-source projects?
Our study compares LLM-generated code summaries against human-written ones collected from open-source projects. If LLMs surpass the quality of open-source code summaries, using them as references is undesirable. Instead, curated code summaries or reference-free evaluation techniques would be necessary. However, even as language model performance improves, calibration remains important; this paper clarifies, and lays out methodologies for studying calibration, even under future improved summarization and evaluation methods. It should be noted, however, that it’s not yet clear to what degree LLM-generated summaries surpass open-source ones in helping developers. Existing work suggests LLMs may surpass humans in writing code summaries (Sun et al., 2024; Su et al., 2024). (Su and McMillan, 2024) find that participants generally preferred GPT-3.5 generated code summaries over open-source ones, often rating them higher in terms of completeness and accuracy. However, LLM-generated summaries and LLM evaluations have not been explicitly judged based on on-task programming performance; the evidence that LLM-generated summaries are better than human summaries relies on human ratings, rather than help with on-task performance; (Stapleton et al., 2020) find that human ratings are uncorrelated with performance on comprehension questions, thus casting doubt on human ratings of completeness and accuracy with respect to code. It’s important to note here that the human ratings we use are for the simpler purpose of selecting a threshold to evaluate the similarity to available human-written summaries (for our subsequent experiments), not to evaluate completeness and accuracy with respect to the code. Our use does not require the human ratings to be based on the same kind of working knowledge of the codebase and software process that the original human summary writer would have possessed.
@@ 6.2. Limitation: Summary Intent Capture
Human developers have different intents (how/what/why) when writing summaries (Hu et al., 2022; Geng et al., 2024), so a user might request summaries to be generated for a particular intent (Su et al., 2024). However, these different intents are not explicitly captured by similarity metrics. A reliable way of integrating intent into automated evaluation metrics and confidence scores remains for future work.
@@ 6.3. Limitation: Platt scaling
Platt scaling is a popular and widely used rescaling strategy. However, another very popular rescaling technique is temperature scaling, which is found to be better than Platt scaling in certain scenarios. In this paper, we could not use temperature scaling because it requires access to complete softmax layers, which are unavailable from the OpenAI [^3] https://openai.com and TogetherAI [^4] https://www.together.ai APIs. In the future, we hope to incorporate temperature when we have access to the complete softmax layer.
@@ 6.4. Self-Reflection Calibration
By “reflective,” we mean the ability of an instruction-tuned model to evaluate its own output. We used two reflective measures to compute model confidence: a logit-based one and a verbalized one.
For the logit-based measure, we presented the LLM with the code and model-generated summary, then asked it whether the summary resembled something a human might write by generating a True/False response. Our prompts were adapted from Spiess et al. (Spiess et al., 2025). Using API access to the top-5 token probabilities, we extracted the logit corresponding to the token True as the model’s confidence score.
The logit-based measure was only feasible with GPT-3.5-Turbo. Other models, like CodeLlama-70B, failed to generate valid Boolean outputs. Additionally, due to API limitations, we didn’t have access to the top-5 log-probabilities for models other than GPT-3.5-Turbo. GPT-3.5-Turbo exhibited severe overconfidence: 94% of predictions had confidence scores above 0.9 and its performance yielded a negative skill score. Rescaling the logits transformed them to near the unskilled base rate.
For the verbalized measure, we asked models to rate the summary’s resemblance to what a human might write on a scale from 1 to 4. We evaluated both zero-shot and few-shot prompting; our prompts followed the human evaluation framework of  (Haque et al., 2022). In the zero-shot setting, CodeLlama-70B failed to produce valid scores, but it succeeded under few-shot prompting using 4 examples from the Haque dataset. Still, in both prompting setups, model outputs were poorly calibrated. We experimented with higher temperatures and majority voting across multiple samples, but these techniques did not improve calibration.
@@ 7. Related Work
8. Conclusion & Acknowledgements
Code summaries are valuable to maintainers, and LLMs can generate good summaries from code. However, LLMs generate summaries that vary in quality: sometimes they resemble human-written summaries, and sometimes not. This indicates a need for developers to know when a summary might be sufficiently similar to a human-written summary. This is a calibration problem: can the model provide a reliable indication of its confidence that a generated summary is “good enough”? We address this problem by selecting a metric known from the literature to be a good proxy for similarity to human-produced summaries, and then thresholding this metric to get an indication of sufficient similarity. We then evaluate the calibration of a range of prompting approaches, together with Platt scaling. We find that in some cases, the first few tokens may provide the best calibration; we also find both retrieval-augmented few-shot prompting and ASAP can be advantageous. We also find that reflection-based approaches do not provide better calibration. To our knowledge, we are the first to evaluate whether LLM-based approaches to code summarization are well calibrated; our work contributes both a methodology, and several useful approaches, to this important question.
We gratefully acknowledge support from NSF CISE, Grant Number 2107592, entitled: “SHF:Medium: Studying and Exploiting the Bimodality of Software”.
@@ 9. Data Availability
Our code and datasets are made available on GitHub.